Practice Exercise 11-1 Pharoah Company purchases equipment on January 1, Year 1, at a cost of $267,000. The asset is expected to have a service life of 6 years and a salvage value of $20,000. Compute the amount of depreciation for each of Years 1 and 2 using the straight-line depreciation method. Depreciation for Year1 Depreciation for Year 2 s Compute the amount of depreciation for each of Years 1 and 2 using the sum-of-the-years-digits method Depreciation for Year 1 Depreciation for Year 2 Compute the amount of depreciation for each of Years 1 and 2 using the double-declining-balance method. (Round answers to O decimal places, e.g. 45,892.) Depreciation for Year 1 Depreciation for Year 2 s
Answer #1

Answers

  • All working forms part of the answer
  • Straight Line Method

A

Cost

$ 2,67,000.00

B

Residual Value

$      20,000.00

C=A - B

Depreciable base

$ 2,47,000.00

D

Life [in years]

6

E=C/D

Annual SLM depreciation

$      41,166.67

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$     2,67,000.00

$     41,166.67

$          2,25,833.33

$             41,166.67

2

$     2,25,833.33

$     41,166.67

$          1,84,666.67

$             82,333.33

  • Sum of Digit Method

A

Cost

$ 2,67,000.00

B

Residual Value

$      20,000.00

C=A – B

Depreciable base

$ 2,47,000.00

1+2+3+4+5+6

Sum of digits

21

Year

Depreciable base

Formula

Depreciation expense

Accumulated Depreciation

Ending Book Value

1

$     2,47,000.00

[247000 x 6/21]

$             70,571.43 or $70,571

$             70,571.43

$ 1,96,428.57

2

$     2,47,000.00

[247000 x 5/21]

$             58,809.52 or $58,810

$          1,29,380.95

$ 1,37,619.05

  • Double Declining Balance Method

A

Cost

$ 2,67,000.00

B

Residual Value

$      20,000.00

C=A - B

Depreciable base

$ 2,47,000.00

D

Life [in years]

6

E=C/D

Annual SLM depreciation

$      41,166.67

F=E/C

SLM Rate

16.67%

G=F x 2

DDB Rate

33.33%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

1

$     2,67,000.00

33.33%

$89,000

2

$     1,78,000.00

33.33%

$59,333